### v2

# TinyIO: Lightweight Reparameterized Inertial Odometry

Abstract
Inertial odometry (IO) is a widely used approach for localization on mobile devices; however, obtaining a lightweight IO model that also achieves high accuracy remains challenging. To address this issue, we propose TinyIO, a lightweight IO method. During training, we adopt a multi-branch architecture to extract diverse motion features more effectively. At inference time, the trained multi-branch model is converted into an equivalent single-path architecture to reduce computational complexity. We further introduce a sparse gating unit that suppresses noise via element-wise selection and integrates contextual motion cues with local, fine-grained features. Extensive experiments on public benchmarks demonstrate that our method attains a favorable trade-off between accuracy and model size. On the RoNIN dataset, TinyIO reduces the ATE by 2.59% compared with RoNIN ResNet and decreases the parameter count by 3.86%.
Index Terms—  Reparameterization, Lightweight, Inertial Measurement Unit signals, Inertial Odometry
1 Introduction
Inertial odometry (IO) aims to estimate a platform’s position over time from inertial measurement unit (IMU) signals[1]. An IMU typically comprises accelerometers and gyroscopes, which measure linear acceleration and angular velocity, respectively. Owing to their low power consumption and inherent protection of data privacy, IO has been widely adopted in assistive localization for individuals with disabilities, positioning of IoT devices[2, 3], and ambient-assisted living [4, 5, 6].
IO methods can generally be categorized into two types: Newtonian-mechanics-based approaches and data-driven approaches. Traditional methods grounded in Newtonian mechanics offer the advantage of low computational overhead; however, they suffer from cumulative errors, and are often constrained by environment-specific factors, which limit their practical applicability[7, 8, 9].
In contrast, data-driven IO methods learn motion patterns directly from IMU signals to estimate position. These methods typically achieve higher localization accuracy and exhibit improved generalization across diverse environments [10, 11, 5, 12, 13]. Nevertheless, many existing approaches emphasize accuracy by employing large-scale models while overlooking the importance of lightweight models for IO, particularly for mobile deployment [2, 3].
To address this issue, we propose a lightweight IO framework named TinyIO, which decouples the training and inference phases to enhance parameter efficiency. The main contributions of this paper are as follows:
  • We perform inference-time reparameterization to decouple the model from its training architecture, thereby improving parameter efficiency.
  • We introduce a sparse gating unit that suppresses noise in IMU signals while effectively fusing contextual motion information with local fine-grained features.
  • Experimental results on multiple public datasets demonstrate that TinyIO achieves a favorable trade-off between accuracy and parameters, as illustrated in Fig. 1.
Fig. 1: Comparison of ATE performance across different models on the RNIN dataset. Each point corresponds to an IO algorithm; points closer to the bottom-left corner indicate lower trajectory error with fewer model parameters. The proposed methods, Tiny and Tiny_deploy, achieve the lowest ATE while maintaining a compact parameter size.
Fig. 2: Schematic diagram of the proposed TinyIO architecture.
2 Method
2.1 Overall Pipeline
The overall pipeline of the proposed TinyIO, illustrated in Fig. 2(a), employs a symmetric architecture. The TinyIO backbone comprises seven stacked Lightweight Inertial Odometry Blocks (LIOBs), with two Reparameterization Blocks (RepBlocks) placed at the input and output for channel alignment. These modules are arranged into five stages to extract rich motion information.
Each LIOB contains a RepBlock and a Sparse Gated Unit (SGU). The RepBlock adopts a multi-branch structure to capture complex motion patterns during training. At inference time, this multi-branch structure is reparameterized into an equivalent single-path feedforward form to improve parameter efficiency, as shown in Fig. 2(b). Feature fusion is performed by the SGU, which applies sparse attention[14, 15] to suppress noise by selectively filtering elements, and fuses contextual motion information with local fine-grained features via a gating mechanism (see Fig. 2(d)).
Finally, a multi-layer perceptron (MLP) head predicts velocity; the network is trained by minimizing the mean squared error (MSE) loss.
2.2 Lightweight Inertial Odometry Block
To balance model complexity and accuracy, we incorporate reparameterization techniques into the Lightweight Inertial Odometry Blocks (LIOBs) and introduce a sparse gating mechanism. Given an IMU signal $\mathbf{X}\in\mathbb{R}^{C\times F}$, where $C=6$ denotes the six channels from the three-axis accelerometer and gyroscope and $F$ denotes the number of frames per time window (i.e., the IMU sampling rate), the processing within each LIOB is:
$\mathbf{X}^{\prime}$ $=\mathrm{RepBlock}(\mathbf{X}),$ (1)
$\mathbf{X}^{\prime\prime}$ $=\mathrm{SGU}(\mathbf{X}^{\prime}).$
Here, $\mathbf{X}^{\prime}$ and $\mathbf{X}^{\prime\prime}$ denote the outputs of the RepBlock and the SGU, respectively.
Fig. 3: Comparison of the basic modules of LIOB and ResNet.
Reparameterization Block (RepBlock). The RepBlock is a core component of each LIOB and is also applied at both the beginning and the end of the backbone for channel alignment. To improve parameter efficiency and enhance feature extraction, the RepBlock decouples the training and inference architectures [16]. Fig. 3 illustrates the structural differences between the RepBlock and the core block used in the RoNIN ResNet backbone.
During training, a multi-branch architecture—similar to that of ResNet—is adopted to extract features from the IMU signal. Specifically, the computation within each RepBlock during training is given by
$\mathbf{Y}=\mathbf{X}+\phi(\mathbf{X})+\psi(\mathbf{X})\in\mathbb{R}^{C\times F},$ (2)
where $\mathbf{X}$ denotes the input to the block, and $\phi(\cdot)$ and $\psi(\cdot)$ denote 1D convolutional branches with kernel sizes 3 and 1, respectively.
At inference, the multi-branch structure is reparameterized into a single-path connection, as shown in Fig. 3(c). This transformation preserves the equivalence of the learned parameters while simplifying the architecture, thereby improving computational efficiency without degrading predictive accuracy. At inference time the operation is represented as
$\mathbf{Y}=\gamma(\mathbf{X})\in\mathbb{R}^{C\times F},$ (3)
where $\gamma(\cdot)$ denotes a single 1D convolution (kernel size 3) obtained by merging the parameters of $\phi(\cdot)$, $\psi(\cdot)$, and the residual (identity) connection from the training phase.
Table 1: Comparison of trajectory prediction performance, ablation studies, and model efficiency. For comparison (left of dashed line), the best results are highlighted in bold. For ablation studies (right of dashed line), the best results are underlined. Parameters and FLOPs: the best results are highlighted in bold. ATE and RTE are measured in meters. Parameters and FLOPs are reported in millions.
Dataset Metric RoNIN ResNet RoNIN LSTM RNIN IMUNet TLIO TinyIO_ deploy TinyIO TinyIO vs RoNIN ResNet Ablation Study
TinyIO (w/o SGU) Full TinyIO
RoNIN ATE 4.043 6.070 7.352 4.093 11.778 4.081 3.939 2.59% 4.288 3.939
RTE 3.123 3.588 3.420 2.859 3.856 2.806 2.857 8.50% 3.175 2.857
RIDI ATE 1.987 5.208 2.259 2.180 2.347 1.980 1.718 13.54% 1.860 1.718
RTE 2.125 5.540 2.426 2.366 2.291 2.123 1.927 9.33% 1.990 1.927
RNIN ATE 2.089 3.057 2.364 2.010 1.970 1.876 1.875 10.25% 2.085 1.875
RTE 1.819 3.264 2.536 1.876 1.785 1.732 1.731 4.82% 1.802 1.731
TLIO ATE 1.601 2.144 1.653 1.609 2.442 1.617 1.459 8.85% 1.560 1.459
RTE 3.968 5.263 4.132 4.113 6.799 4.135 3.637 8.33% 3.916 3.637
FLOPs (M) 38.252 54.023 72.032 18.838 39.436 49.379 60.968 - -
Param (M) 4.635 10.719 5.259 3.659 5.424 3.809 4.459 - -
Sparse Gated Unit (SGU). While the RepBlock emphasizes local fine-grained motion features, it lacks sufficient contextual awareness. To address this limitation, we introduce the SGU, which suppresses noise through element-wise filtering and integrates contextual motion information with local details to enable multi-scale feature interaction. The SGU structure is illustrated in Fig. 2(d). It comprises two branches: a value branch that captures contextual dependencies via sparse attention, and a gating branch that captures local fine-grained variations.
In the value branch, we apply a standard 1D convolution and a depthwise 1D convolution (followed by a transpose) to produce keys $K$, queries $Q$, and values $V$, where $K,Q,V\in\mathbb{R}^{F\times C}$. We then compute the attention score matrix $\frac{QK^{\top}}{\sqrt{C}}\in\mathbb{R}^{F\times F}$, which captures pairwise dependencies among frames. To reduce spurious dependencies and suppress noise[15], we apply a max-$e$ sparsification that retains only the top $e$ fraction of elements in each row. (In our experiments we set $e=0.8$, i.e., retain the top 80% of elements per row.) The sparse score matrix is normalized row-wise by a softmax and then applied to $V$ to obtain the attention output:
$\mathrm{Value}\;=\;\mathrm{Softmax}\!\left(\mathrm{max}\text{-}e\!\left(\frac{K^{\top}Q}{\sqrt{C}}\right)\right)V\;\in\;\mathbb{R}^{F\times C}.$ (4)
The gating branch applies a depthwise convolution followed by a nonlinear activation to capture local, channel-wise motion variations. This branch functions equivalently to a gated channel-attention module and increases sensitivity to subtle motion cues that the value branch may overlook.
Finally, the outputs of the value and gating branches are fused via the gating mechanism to facilitate multi-scale feature integration while suppressing implicit noise. The SGU is a versatile module suitable for a range of inertial localization scenarios.
3 Experiments and Analysis
3.1 Experimental Settings
Training details. All training and evaluation procedures are carried out on a single NVIDIA RTX 3090 GPU with 24 GB of memory. The network is optimized using the Adam optimizer with an initial learning rate of $10^{-4}$. Training runs for up to 100 epochs, with early stopping triggered when the learning rate decays below $10^{-6}$. The channel dimensions of TinyIO’s five stages are set to $[64,64,128,256,512]$.
Datasets and metrics. We evaluate on four widely used open-source datasets: RIDI [17], RoNIN [18], RNIN [19], and TLIO [20]. Each dataset is split into training, validation, and test subsets with a ratio of 8:1:1, and all models are retrained on the respective datasets for fair comparison. Three evaluation metrics are used: absolute trajectory error (ATE) [21], relative trajectory error (RTE) [21], and position drift error (PDE) [22]. ATE quantifies global trajectory consistency, RTE measures local trajectory drift, and PDE evaluates endpoint deviation.
Baselines. Prior work has shown that data-driven approaches typically outperform Newtonian model-based methods in localization accuracy [17, 23, 19, 20, 24]. Accordingly, we compare TinyIO against representative data-driven baselines: the ResNet and LSTM architectures from RoNIN [18], the network designs of RNIN [19] and TLIO [20], and the lightweight IMUNet architecture [2].
Fig. 4: Visual comparison of TinyIO and RoNIN ResNet on various datasets.
Fig. 5: PDE comparison between TinyIO and three baseline models from RoNIN on the RoNIN dataset.
3.2 Comparisons with the baselines
Metric evaluation. Table 1 (left) summarizes ATE and RTE comparisons between TinyIO and the baseline models. TinyIO and its reparameterized variant, TinyIO_deploy, achieve the lowest localization errors in most scenarios. For example, relative to RoNIN ResNet, TinyIO reduces ATE by 2.59%—13.54% and RTE by 4.82%—9.33% across the four datasets. Model complexity increases progressively from RoNIN to TLIO and RNIN, which raises deployment challenges for edge devices. Although IMUNet uses depthwise-separable convolutions to reduce complexity, its localization accuracy remains limited. In contrast, TinyIO decouples training and inference to substantially reduce parameter counts, while the SGU suppresses implicit noise and fuses local fine-grained and contextual motion features to strike an effective balance between accuracy and compactness. Overall, these improvements explain the superior performance of TinyIO over the baselines.
Visualization. Fig. 5 shows a comparison of PDE among TinyIO, TinyIO_deploy, and RoNIN (results on the RoNIN dataset are shown due to space constraints). TinyIO attains the smallest PDE median and the most compact error distribution, indicating superior endpoint approximation. Although reparameterization slightly increases PDE for TinyIO_deploy, TinyIO_deploy still outperforms RoNIN ResNet and preserves high-fidelity trajectory reconstruction consistent with the training-phase model.
Fig. 4 visualizes trajectory reconstructions produced by TinyIO and the baseline models on the test sets. In Fig. 4(d), TinyIO effectively captures contextual motion information, yielding accurate and stable tracking. After reparameterization, the positioning error of TinyIO_deploy increases slightly but remains lower than that of RoNIN ResNet, suggesting that the accuracy degradation due to reparameterization is within acceptable bounds. For complex trajectories (Fig. 4(a)–(c)), TinyIO better models local fine-grained motion and exhibits substantially less drift than RoNIN ResNet when handling rotational or highly nonlinear motions; consequently, both TinyIO and TinyIO_deploy closely follow ground-truth trajectories, whereas RoNIN ResNet shows pronounced drift during turns and increasingly larger deviations as trajectory complexity grows.
3.3 Ablation study
We conduct ablation experiments to validate the reparameterization strategy and the contributions of individual modules (SGU and RepBlock).
The middle section of Table 1 reports performance comparisons between TinyIO and TinyIO_deploy across the four datasets. After reparameterization, TinyIO_deploy (3.809M parameters) reduces the parameter count by approximately 14.52% relative to TinyIO (4.456M), and its model size is only 82.18% that of RoNIN ResNet (4.635M). Despite this compression, the increases in ATE and RTE for TinyIO_deploy are marginal, and TinyIO_deploy still outperforms RoNIN ResNet in localization accuracy. Visualization results in Fig. 4 corroborate these findings. These results indicate that the proposed reparameterization achieves a favorable trade-off between parameters and accuracy, making TinyIO particularly suitable for edge devices with limited storage.
The right portion of Table 1 shows module ablation results. Removing SGU (TinyIO w/o SGU) leaves only the RepBlock: the model still captures local fine-grained motion and yields substantially lower localization error than RoNIN ResNet, indicating that RepBlock both reduces parameter complexity and contributes positively to accuracy. Integrating SGU into the full TinyIO architecture further improves accuracy, demonstrating that multi-scale feature fusion via SGU is critical for enhanced trajectory estimation. These findings confirm the effectiveness and complementarity of the proposed components.
4 Conclusion
We present TinyIO, a lightweight IO framework that attains a favorable trade-off between localization accuracy and model parameters. In practical deployment, however, additional procedures such as model fine-tuning may be necessary to adapt to previously unseen carrier motion patterns.
5 Acknowledgements
This work is supported by Science and Technology Major Program of Fujian Province (No. 2022HZ026007) and partly supported by the Science and Technology Planning Project of Fujian province (2022I0001).

### v4

@@
 At inference time, the trained multi-branch model is converted into an equivalent single-path architecture to reduce computational complexity.
-We further introduce a sparse gating unit that suppresses noise via element-wise selection and integrates contextual motion cues with local, fine-grained features.
-Extensive experiments on public benchmarks demonstrate that our method attains a favorable trade-off between accuracy and model size.
-On the RoNIN dataset, TinyIO reduces the ATE by 2.59% compared with RoNIN ResNet and decreases the parameter count by 3.86%.
-Index Terms—  Reparameterization, Lightweight, Inertial Measurement Unit signals, Inertial Odometry
-1 Introduction
-Inertial odometry (IO) aims to estimate a platform’s position over time from inertial measurement unit (IMU) signals[1].
+We further propose a Dual-Path Adaptive Attention mechanism (DPAA), which enhances TinyIO’s perception of contextual motion along both channel and temporal dimensions with negligible additional parameters.
+Extensive experiments on public datasets demonstrate that our method attains a favorable trade-off between accuracy and model size.
+On the RoNIN dataset, TinyIO reduces the ATE by 23.53% compared with R-ResNet and decreases the parameter count by 3.68%.
+Code repository will be made available upon completion of the double-blind review process.
+I INTRODUCTION
+Inertial odometry (IO) aims to estimate a platform’s position over time from inertial measurement unit (IMU) signals [6, 15].
 An IMU typically comprises accelerometers and gyroscopes, which measure linear acceleration and angular velocity, respectively.
-Owing to their low power consumption and inherent protection of data privacy, IO has been widely adopted in assistive localization for individuals with disabilities, positioning of IoT devices[2, 3], and ambient-assisted living [4, 5, 6].
+Owing to their low power consumption, independence from external infrastructure, and inherent data privacy preservation, IO has been widely adopted in assistive localization for individuals with disabilities, indoor positioning, and ambient-assisted living [7, 29, 21, 22, 30, 16].
 IO methods can generally be categorized into two types: Newtonian-mechanics-based approaches and data-driven approaches.
-Traditional methods grounded in Newtonian mechanics offer the advantage of low computational overhead; however, they suffer from cumulative errors, and are often constrained by environment-specific factors, which limit their practical applicability[7, 8, 9].
+Traditional methods grounded in Newtonian mechanics offer the advantage of low computational overhead; however, they suffer from cumulative errors and are often constrained by environment-specific factors, which limit their practical applicability [18, 10, 11].
 In contrast, data-driven IO methods learn motion patterns directly from IMU signals to estimate position.
 These methods typically achieve higher localization accuracy and exhibit improved generalization across diverse environments [30, 16, 1, 23, 9].
-Nevertheless, many existing approaches emphasize accuracy by employing large-scale models while overlooking the importance of lightweight models for IO, particularly for mobile deployment [2, 3].
+Nevertheless, many existing approaches prioritize accuracy by employing large-scale models while neglecting the need for lightweight architectures suitable for mobile deployment [29, 21].
 To address this issue, we propose a lightweight IO framework named TinyIO, which decouples the training and inference phases to enhance parameter efficiency.
@@
   • We perform inference-time reparameterization to decouple the model from its training architecture, thereby improving parameter efficiency.
-  • We introduce a sparse gating unit that suppresses noise in IMU signals while effectively fusing contextual motion information with local fine-grained features.
-  • Experimental results on multiple public datasets demonstrate that TinyIO achieves a favorable trade-off between accuracy and parameters, as illustrated in Fig. 1.
-Fig.
-1: Comparison of ATE performance across different models on the RNIN dataset.
+  • We propose a Dual-Path Adaptive Attention mechanism (DPAA) that enhances the network’s modeling capability of contextual motion information along both channel and temporal dimensions, introducing only 0.037 M additional parameters.
+  • Experimental results on multiple public datasets demonstrate that TinyIO achieves a favorable trade-off between accuracy and model size, as illustrated in Fig. 1.
+Figure 1: Comparison of ATE performance across different models on the OxIOD dataset.
 Each point corresponds to an IO algorithm; points closer to the bottom-left corner indicate lower trajectory error with fewer model parameters.
-The proposed methods, Tiny and Tiny_deploy, achieve the lowest ATE while maintaining a compact parameter size.
-Fig. 2: Schematic diagram of the proposed TinyIO architecture.
-2 Method
-2.1 Overall Pipeline
-The overall pipeline of the proposed TinyIO, illustrated in Fig.
-2(a), employs a symmetric architecture.
-The TinyIO backbone comprises seven stacked Lightweight Inertial Odometry Blocks (LIOBs), with two Reparameterization Blocks (RepBlocks) placed at the input and output for channel alignment.
-These modules are arranged into five stages to extract rich motion information.
-Each LIOB contains a RepBlock and a Sparse Gated Unit (SGU).
-The RepBlock adopts a multi-branch structure to capture complex motion patterns during training.
+Figure 2: The overall architecture of the proposed TinyIO, comprising the Reparameterization Block (RepBlock) and the Dual-path Adaptive Attention mechanism (DPAA).
+II RELATED WORK
+II-A Newtonian-mechanics-based approaches
+Traditional inertial localization methods are typically based on Newtonian mechanics.
+For example, SINS [18] rotates the acceleration vector into a global coordinate system to compensate for gravity and then estimates displacement through double integration.
+However, measurement noise continuously accumulates during this process, leading to progressive drift in the estimated trajectory.
+To mitigate this issue, researchers have introduced physical priors to correct errors.
+For instance, PDR [13] exploits the periodicity of human motion to detect gait and iteratively updates the trajectory using estimated step length and heading.
+ZUPT [20], utilizing foot-mounted sensors, applies a zero-velocity constraint during stance phases to correct accumulated errors and improve accuracy under stationary conditions.
+Although these methods effectively suppress error propagation, their reliance on specific physical priors limits their applicability to a narrow range of motion states and scenarios, reducing robustness in diverse environments.
+II-B Data-driven approaches
+Data-driven methods have demonstrated superior adaptability across diverse application scenarios, exhibiting greater robustness and improved accuracy compared to traditional approaches[19, 27].
+For example, IONet [2] and RoNIN [8] disregard the IMU’s mounting position and directly infer motion from raw measurements using a unified network framework, irrespective of how the device is worn.
+Specifically, RoNIN employs ResNet as the inference backbone; its robust localization performance and straightforward architecture have laid a solid foundation for subsequent research.
+Building on this, TLIO [12] and LIDR [28] integrate the SCEKF and the Invariant Extended Kalman Filter (LIEKF), respectively, to further refine ResNet’s outputs and mitigate localization errors.
+RNIN-VIO [4] enhances the modeling of long-term dependencies by combining LSTM with R-ResNet.
+IMUNet[29], DeepILS[21], and Deeplite[22] attempt to adopt lightweight architectures to improve parameter efficiency but achieve limited localization accuracy.
+In contrast, DIO contends that pose variations resulting from device placement are significant and introduces a novel pose-invariant inertial tracking loss, which regresses both the velocity magnitude and the trigonometric representation of the forward direction [25].
+SCHNN [24] and SSHNN [26] extend DIO by incorporating CNN and LSTM architectures augmented with CNN-based attention mechanisms.
+Similarly, RIO [1] and EqNIO [9] enrich the inference network with components designed to exploit rotational equivariance.
+More recently, CTIN [17] and iMOT [14] have explored the use of Transformer architectures for inertial localization, aiming to capture contextual dependencies, and have achieved promising results.
+While the adoption of increasingly complex network architectures has enhanced accuracy, it often comes at the cost of parameter efficiency, posing significant challenges for deployment on resource-constrained devices.
+To address this, we propose a lightweight localization algorithm that achieves a balance between accuracy and parameter efficiency by incorporating a lightweight attention mechanism and decoupling the training and inference phases.
+III METHOD
+III-A Overall Pipeline
+The overall pipeline of the proposed TinyIO is illustrated in Fig.
+2.
+The TinyIO backbone primarily consists of Lightweight Inertial Odometry Blocks (LIOBs), which infer human motion velocity from acceleration and angular velocity signals.
+Each LIOB contains a Reparameterization Blocks (RepBlocks) and a Dual-Path Adaptive Attention mechanism (DPAA).
+The RepBlock adopts a multi-branch structure during training to capture complex motion patterns.
 At inference time, this multi-branch structure is reparameterized into an equivalent single-path feedforward form to improve parameter efficiency, as shown in Fig.
 2(b).
-Feature fusion is performed by the SGU, which applies sparse attention[14, 15] to suppress noise by selectively filtering elements, and fuses contextual motion information with local fine-grained features via a gating mechanism (see Fig.
-2(d)).
-Finally, a multi-layer perceptron (MLP) head predicts velocity; the network is trained by minimizing the mean squared error (MSE) loss.
-2.2 Lightweight Inertial Odometry Block
-To balance model complexity and accuracy, we incorporate reparameterization techniques into the Lightweight Inertial Odometry Blocks (LIOBs) and introduce a sparse gating mechanism.
+The DPAA compensates for the RepBlock’s limited ability to model long-range dependencies by adaptively aggregating motion context along both the temporal and channel dimensions, adhering to the principle of parameter efficiency and introducing only 0.037M additional parameters.
+Finally, a multi-layer perceptron (MLP) head predicts velocity, and the network is trained by minimizing the mean squared error (MSE) loss.
+Figure 3: Comparison of the basic modules of LIOB and ResNet.
+III-B Lightweight Inertial Odometry Block
+To balance model complexity and accuracy, we incorporate reparameterization techniques into the Lightweight Inertial Odometry Blocks (LIOBs) and propose a dual-path adaptive attention mechanism.
 Given an IMU signal $\mathbf{X}\in\mathbb{R}^{C\times F}$, where $C=6$ denotes the six channels from the three-axis accelerometer and gyroscope and $F$ denotes the number of frames per time window (i.e., the IMU sampling rate), the processing within each LIOB is:
-$\mathbf{X}^{\prime}$ $=\mathrm{RepBlock}(\mathbf{X}),$ (1)
-$\mathbf{X}^{\prime\prime}$ $=\mathrm{SGU}(\mathbf{X}^{\prime}).$
-Here, $\mathbf{X}^{\prime}$ and $\mathbf{X}^{\prime\prime}$ denote the outputs of the RepBlock and the SGU, respectively.
-Fig. 3: Comparison of the basic modules of LIOB and ResNet.
+$\mathbf{X}^{\prime}$ $=\mathrm{RepBlock}(\mathbf{X})$ (1)
+$\mathbf{X}^{\prime\prime}$ $=\mathrm{DPAA}(\mathbf{X}^{\prime})$
+Here, $\mathbf{X}^{\prime}$ and $\mathbf{X}^{\prime\prime}$ denote the outputs of the RepBlock and the DPAA, respectively.
 Reparameterization Block (RepBlock).
@@
 Fig.
-3 illustrates the structural differences between the RepBlock and the core block used in the RoNIN ResNet backbone.
+3 illustrates the structural differences between the RepBlock and the core block used in the R-ResNet[8] backbone.
 During training, a multi-branch architecture—similar to that of ResNet—is adopted to extract features from the IMU signal.
 Specifically, the computation within each RepBlock during training is given by
-$\mathbf{Y}=\mathbf{X}+\phi(\mathbf{X})+\psi(\mathbf{X})\in\mathbb{R}^{C\times F},$ (2)
+$\mathbf{X}^{\prime}=\mathbf{X}+\phi(\mathbf{X})+\psi(\mathbf{X})\in\mathbb{R}^{C\times F}$ (2)
 where $\mathbf{X}$ denotes the input to the block, and $\phi(\cdot)$ and $\psi(\cdot)$ denote 1D convolutional branches with kernel sizes 3 and 1, respectively.
@@
 3(c).
-This transformation preserves the equivalence of the learned parameters while simplifying the architecture, thereby improving computational efficiency without degrading predictive accuracy.
-At inference time the operation is represented as
-$\mathbf{Y}=\gamma(\mathbf{X})\in\mathbb{R}^{C\times F},$ (3)
+This transformation preserves the functional equivalence of the learned representation while simplifying the architecture, thereby improving computational efficiency without degrading predictive accuracy.
+At inference time, the operation is represented as
+$\mathbf{X}^{\prime}=\gamma(\mathbf{X})\in\mathbb{R}^{C\times F}$ (3)
 where $\gamma(\cdot)$ denotes a single 1D convolution (kernel size 3) obtained by merging the parameters of $\phi(\cdot)$, $\psi(\cdot)$, and the residual (identity) connection from the training phase.
-Table 1: Comparison of trajectory prediction performance, ablation studies, and model efficiency.
-For comparison (left of dashed line), the best results are highlighted in bold.
-For ablation studies (right of dashed line), the best results are underlined.
-Parameters and FLOPs: the best results are highlighted in bold.
+Dual-Path Adaptive Attention (DPAA).
+While the RepBlock emphasizes local fine-grained motion features, it lacks sufficient contextual awareness.
+To address this limitation, we propose the DPAA, which models contextual motion information along both the channel and temporal dimensions through a lightweight attention mechanism and then adaptively fuses them to achieve a comprehensive understanding of the IMU signal.
+The DPAA structure is illustrated in Fig.
+2(d).
+It comprises two branches: a lightweight channel attention branch that models inter-channel dependencies, and a lightweight temporal attention branch that captures the importance of different time steps.
+In the lightweight channel attention branch, we apply adaptive average pooling and adaptive max pooling to the output feature $\mathbf{X}^{\prime}$ from the RepBlock to extract channel-wise statistics.
+These statistics are then fused through a two-layer convolutional module with shared parameters and mapped to channel weights in $[0,1]$ via a sigmoid function.
+The resulting channel weights modulate $\mathbf{X}^{\prime}$ to emphasize informative channels.
+Formally, this process is expressed as:
+$\mathbf{z}_{\text{avg}}$ $=\frac{1}{F}\sum_{f=1}^{F}\mathbf{X}^{\prime}_{:,f}\in\mathbb{R}^{C}$ (4)
+$\mathbf{z}_{\text{max}}$ $=\max_{f\in\{1,\dots,F\}}\mathbf{X}^{\prime}_{:,f}\in\mathbb{R}^{C}$ (5)
+$\mathbf{w}_{c}$ $=\sigma\Big(\mathbf{W}_{2}\,\delta(\mathbf{W}_{1}\mathbf{z}_{\text{avg}})+\mathbf{W}_{2}\,\delta(\mathbf{W}_{1}\mathbf{z}_{\text{max}})\Big)\in\mathbb{R}^{C}$ (6)
+$\mathbf{Y}_{\text{ca}}$ $=\mathbf{X}^{\prime}\odot\mathbf{w}_{c}^{\top}\in\mathbb{R}^{C\times F}$ (7)
+where $\mathbf{z}_{\text{avg}}$ and $\mathbf{z}_{\text{max}}$ denote the channel statistics obtained via adaptive average and max pooling, respectively; $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are learnable 1D convolutional weights (effectively functioning as shared fully connected layers); $\delta$ and $\sigma$ represent the ReLU and sigmoid activation functions, respectively; $\mathbf{w}_{c}$ is the fused channel weight vector; and $\odot$ denotes broadcasting element-wise multiplication.
+Notably, the use of shared-weight convolutions instead of standard fully connected layers further reduces parameter count, satisfying the constraints of resource-constrained deployment.
+TABLE I: Comparison of trajectory prediction performance and model efficiency.
 ATE and RTE are measured in meters.
-Parameters and FLOPs are reported in millions.
-Dataset Metric RoNIN ResNet RoNIN LSTM RNIN IMUNet TLIO TinyIO_ deploy TinyIO TinyIO vs RoNIN ResNet Ablation Study
-TinyIO (w/o SGU) Full TinyIO
-RoNIN ATE 4.043 6.070 7.352 4.093 11.778 4.081 3.939 2.59% 4.288 3.939
-RTE 3.123 3.588 3.420 2.859 3.856 2.806 2.857 8.50% 3.175 2.857
-RIDI ATE 1.987 5.208 2.259 2.180 2.347 1.980 1.718 13.54% 1.860 1.718
-RTE 2.125 5.540 2.426 2.366 2.291 2.123 1.927 9.33% 1.990 1.927
-RNIN ATE 2.089 3.057 2.364 2.010 1.970 1.876 1.875 10.25% 2.085 1.875
-RTE 1.819 3.264 2.536 1.876 1.785 1.732 1.731 4.82% 1.802 1.731
-TLIO ATE 1.601 2.144 1.653 1.609 2.442 1.617 1.459 8.85% 1.560 1.459
-RTE 3.968 5.263 4.132 4.113 6.799 4.135 3.637 8.33% 3.916 3.637
-FLOPs (M) 38.252 54.023 72.032 18.838 39.436 49.379 60.968 - -
-Param (M) 4.635 10.719 5.259 3.659 5.424 3.809 4.459 - -
-Sparse Gated Unit (SGU).
-While the RepBlock emphasizes local fine-grained motion features, it lacks sufficient contextual awareness.
-To address this limitation, we introduce the SGU, which suppresses noise through element-wise filtering and integrates contextual motion information with local details to enable multi-scale feature interaction.
-The SGU structure is illustrated in Fig.
-2(d).
-It comprises two branches: a value branch that captures contextual dependencies via sparse attention, and a gating branch that captures local fine-grained variations.
-In the value branch, we apply a standard 1D convolution and a depthwise 1D convolution (followed by a transpose) to produce keys $K$, queries $Q$, and values $V$, where $K,Q,V\in\mathbb{R}^{F\times C}$.
-We then compute the attention score matrix $\frac{QK^{\top}}{\sqrt{C}}\in\mathbb{R}^{F\times F}$, which captures pairwise dependencies among frames.
-To reduce spurious dependencies and suppress noise[15], we apply a max-$e$ sparsification that retains only the top $e$ fraction of elements in each row.
-(In our experiments we set $e=0.8$, i.e., retain the top 80% of elements per row.) The sparse score matrix is normalized row-wise by a softmax and then applied to $V$ to obtain the attention output:
-$\mathrm{Value}\;=\;\mathrm{Softmax}\!\left(\mathrm{max}\text{-}e\!\left(\frac{K^{\top}Q}{\sqrt{C}}\right)\right)V\;\in\;\mathbb{R}^{F\times C}.$ (4)
-The gating branch applies a depthwise convolution followed by a nonlinear activation to capture local, channel-wise motion variations.
-This branch functions equivalently to a gated channel-attention module and increases sensitivity to subtle motion cues that the value branch may overlook.
-Finally, the outputs of the value and gating branches are fused via the gating mechanism to facilitate multi-scale feature integration while suppressing implicit noise.
-The SGU is a versatile module suitable for a range of inertial localization scenarios.
-3 Experiments and Analysis
-3.1 Experimental Settings
+The best results are highlighted with an orange background, and the second-best results with a green background.
+Dataset Metric R-ResNet R-LSTM RNIN I-IMUNet I-EfficientNet I-MnasNet I-MobileNet DeepILS Deeplite TLIO TinyIO_ deploy TinyIO TinyIO vs R-ResNet
+Publication ICRA 2020 ICRA 2020 ISMAR 2021 TIM 2024 TIM 2024 TIM 2024 TIM 2024 IoT-J 2025 IoT-J 2025 RAL 2020
+RoNIN ATE 6.12 10.38 6.88 6.07 5.84 7.49 4.94 5.80 6.36 5.52 4.55 4.68 23.53%
+RTE 4.08 4.92 4.17 4.04 3.91 4.02 3.87 3.90 4.38 3.80 3.87 3.93 3.68%
+RIDI ATE 2.43 2.82 2.77 3.20 1.93 3.14 2.89 1.89 2.31 2.24 1.89 1.89 22.22%
+RTE 2.60 2.98 2.99 3.24 2.20 3.00 3.00 2.36 2.56 2.52 2.19 2.19 15.77%
+OxIOD ATE 3.50 3.13 3.66 1.64 5.58 3.67 2.85 3.29 2.05 2.02 1.41 1.41 59.71%
+RTE 1.30 1.39 1.34 1.04 1.68 1.46 1.20 1.23 1.11 0.97 0.93 0.93 28.46%
+IMUNet ATE 8.16 12.89 8.40 8.68 6.11 12.23 8.130 7.86 7.46 7.49 6.35 5.66 30.64%
+RTE 4.71 8.71 5.05 5.95 4.60 7.18 5.52 5.59 5.75 5.27 4.44 4.35 7.64%
+FLOPs (M) 38.252 57.473 72.032 18.838 28.831 25.256 34.559 15.287 59.045 39.436 37.595 26.194 31.52%
+Params (M) 4.635 13.589 5.259 3.659 3.232 2.979 3.179 2.291 0.297 5.424 2.975 3.570 22.98%
+Peak Memory (MB) 29.069 80.184 59.159 64.955 61.256 72.833 68.814 70.979 50.526 103.165 22.679 25.609 11.90%
+Figure 4: (a) and (b) show the cumulative distribution function (CDF) curves of the ATE and RTE for four localization algorithms, respectively.
+Since TinyIO and TinyIO_deploy exhibit similar performance, their curves overlap (red and orange curves).
+(c) presents the PDE metric for the four algorithms.
+The lightweight temporal attention branch follows a similar design philosophy.
+We first compute the mean and maximum statistics along the temporal dimension of $\mathbf{X}^{\prime}$, concatenate them into a two-channel temporal feature map, and then apply a 1D convolution (operating along the time axis) to model local temporal dependencies and map the concatenated features to a temporal attention score.
+This score is normalized via sigmoid to produce temporal weights in $[0,1]$, which modulate $\mathbf{X}^{\prime}$ to highlight critical time steps.
+The procedure is formulated as:
+$\mathbf{m}_{\text{avg}}$ $=\frac{1}{C}\sum_{c=1}^{C}\mathbf{X}^{\prime}_{c,:}$ (8)
+$\mathbf{m}_{\text{max}}$ $=\max_{c\in\{1,\dots,C\}}\mathbf{X}^{\prime}_{c,:}$ (9)
+$\mathbf{M}$ $=\begin{bmatrix}\mathbf{m}_{\text{avg}}\\ \mathbf{m}_{\text{max}}\end{bmatrix}$ (10)
+$\mathbf{w}_{t}$ $=\sigma\big(\mathbf{W}_{3}\mathbf{M}\big)$ (11)
+$\mathbf{Y}_{\text{ta}}$ $=\mathbf{X}^{\prime}\odot\mathbf{w}_{t}$ (12)
+where $\mathbf{m}_{\text{avg}}$ and $\mathbf{m}_{\text{max}}$ are the temporal mean and max statistics, $\mathbf{W}_{3}$ is a learnable 1D convolutional kernel applied along the temporal dimension, and $\mathbf{w}_{t}$ is the resulting temporal weight vector.
+Finally, we dynamically fuse $\mathbf{Y}_{\text{ca}}$ and $\mathbf{Y}_{\text{ta}}$ using learnable scalar weights:
+$\mathbf{X}^{\prime\prime}$ $=\alpha\cdot\mathbf{Y}_{\text{ca}}+\beta\cdot\mathbf{Y}_{\text{ta}}$ (13)
+$1$ $=\alpha+\beta$
+where $\alpha$ and $\beta$ are learnable parameters.
+This design enables adaptive, dual-path learning over both channel and temporal dimensions while maintaining high parameter efficiency, making it significantly more suitable for mobile deployment compared to standard attention mechanisms.
+IV EXPERIMENTS AND ANALYSIS
+IV-A Experimental Settings
 Training details.
-All training and evaluation procedures are carried out on a single NVIDIA RTX 3090 GPU with 24 GB of memory.
+All training and evaluation are conducted on an NVIDIA RTX 3090 GPU with 24 GB of memory, using CUDA version 12.2 and PyTorch version 2.5.1.
 The network is optimized using the Adam optimizer with an initial learning rate of $10^{-4}$.
-Training runs for up to 100 epochs, with early stopping triggered when the learning rate decays below $10^{-6}$.
-The channel dimensions of TinyIO’s five stages are set to $[64,64,128,256,512]$.
+Training runs for up to 40 epochs, with early stopping triggered when the learning rate decays below $10^{-6}$.
+The channel dimensions of TinyIO’s four stages are set to $[48,96,192,512]$.
 Datasets and metrics.
-We evaluate on four widely used open-source datasets: RIDI [17], RoNIN [18], RNIN [19], and TLIO [20].
-Each dataset is split into training, validation, and test subsets with a ratio of 8:1:1, and all models are retrained on the respective datasets for fair comparison.
-Three evaluation metrics are used: absolute trajectory error (ATE) [21], relative trajectory error (RTE) [21], and position drift error (PDE) [22].
-ATE quantifies global trajectory consistency, RTE measures local trajectory drift, and PDE evaluates endpoint deviation.
+We evaluate our method on widely used open-source datasets: RIDI [27], RoNIN [8], RNIN [4], and TLIO [12].
+Each dataset is split into training, validation, and test subsets with a ratio of 8:1:1, and all models are retrained on their respective datasets.
+The evaluation metrics include commonly used IO metrics: absolute trajectory error (ATE) [19], relative trajectory error (RTE) [19], and position drift error (PDE) [17].
 Baselines.
-Prior work has shown that data-driven approaches typically outperform Newtonian model-based methods in localization accuracy [17, 23, 19, 20, 24].
-Accordingly, we compare TinyIO against representative data-driven baselines: the ResNet and LSTM architectures from RoNIN [18], the network designs of RNIN [19] and TLIO [20], and the lightweight IMUNet architecture [2].
-Fig. 4: Visual comparison of TinyIO and RoNIN ResNet on various datasets.
-Fig. 5: PDE comparison between TinyIO and three baseline models from RoNIN on the RoNIN dataset.
-3.2 Comparisons with the baselines
+Prior work has demonstrated that data-driven approaches generally outperform Newtonian model-based methods in terms of localization accuracy [27, 3, 4, 12, 14].
+Accordingly, we compare TinyIO against widely used open-source data-driven baselines, including DeepILS [21], Deeplite [22], the R-ResNet and R-LSTM variants from RoNIN [8], the I-IMUNet, I-MobileNet, I-MNasNet, and I-EfficientNet models from IMUNet [29], as well as the architectures proposed in TLIO [12] and RNIN [4].
+Figure 5: Comparison of sampled trajectories on the test set for TinyIO, TinyIO_deploy, and R-ResNet.
+The values in parentheses denote the specific metrics for each algorithm on the corresponding trajectory.
+Due to the similar performance of TinyIO and TinyIO_deploy, their trajectories exhibit overlapping coverage (red and orange curves).
+For the OxIOD dataset, only a segment of the full trajectory is visualized to fit the display scale.
+IV-B Comparisons with the baselines
 Metric evaluation.
-Table 1 (left) summarizes ATE and RTE comparisons between TinyIO and the baseline models.
+Table I summarizes ATE and RTE comparisons between TinyIO and the baseline models.
 TinyIO and its reparameterized variant, TinyIO_deploy, achieve the lowest localization errors in most scenarios.
-For example, relative to RoNIN ResNet, TinyIO reduces ATE by 2.59%—13.54% and RTE by 4.82%—9.33% across the four datasets.
+For example, relative to R-ResNet, TinyIO reduces ATE by 22.22%—59.71% and RTE by 3.68%—28.46% across the all datasets.
 Model complexity increases progressively from RoNIN to TLIO and RNIN, which raises deployment challenges for edge devices.
 Although IMUNet uses depthwise-separable convolutions to reduce complexity, its localization accuracy remains limited.
-In contrast, TinyIO decouples training and inference to substantially reduce parameter counts, while the SGU suppresses implicit noise and fuses local fine-grained and contextual motion features to strike an effective balance between accuracy and compactness.
+In contrast, TinyIO decouples training and inference to substantially reduce parameter counts, while the DPAA models contextual motion information through a lightweight attention mechanism to strike an effective balance between accuracy and compactness.
 Overall, these improvements explain the superior performance of TinyIO over the baselines.
@@
 Fig.
-5 shows a comparison of PDE among TinyIO, TinyIO_deploy, and RoNIN (results on the RoNIN dataset are shown due to space constraints).
+4(a) and (b) present the cumulative distribution function (CDF) curves of the ATE and RTE for TinyIO, TinyIO_deploy, R-LSTM and R-ResNet on the IMUNet dataset, respectively.
+Results on the IMUNet dataset are presented due to space constraints.
+As shown in the figures, the TinyIO curve (orange) is positioned above and to the left of the R-ResNet curve (blue) and R-LSTM (green), indicating that TinyIO achieves higher accuracy compared to R-ResNet and R-LSTM.
+For instance, approximately 82% of trajectories from TinyIO exhibit ATE values below 0.06 m, whereas only approximately 72% of trajectories from R-ResNet achieve ATE values below this threshold.
+Fig.
+4(c) shows a comparison of PDE.
 TinyIO attains the smallest PDE median and the most compact error distribution, indicating superior endpoint approximation.
-Although reparameterization slightly increases PDE for TinyIO_deploy, TinyIO_deploy still outperforms RoNIN ResNet and preserves high-fidelity trajectory reconstruction consistent with the training-phase model.
+Although reparameterization slightly increases the PDE for TinyIO_deploy, it still outperforms R-ResNet and preserves high-fidelity trajectory reconstruction consistent with the training-phase model.
+TABLE II: Ablation study results on open-source datasets. Best results are highlighted in orange, second-best in green.
+Dataset Metric R-ResNet TinyIO (only RepBlock) Full TinyIO (RepBlock&DPAA) Performance (meters) RoNIN ATE 6.12 4.99 4.68 RTE 4.08 3.94 3.93 RIDI ATE 2.43 2.16 1.89 RTE 2.60 2.50 2.19 IMUNet ATE 8.16 7.45 5.66 RTE 4.71 5.13 4.35 OxIOD ATE 3.50 2.28 1.41 RTE 1.30 1.12 0.93 Parameters (M) 4.635 3.533 3.57 FLOPs (M) 38.252 26.019 26.194 Peak Memory (MB) 29.069 25.587 25.609
 Fig.
@@
 In Fig.
-4(d), TinyIO effectively captures contextual motion information, yielding accurate and stable tracking.
-After reparameterization, the positioning error of TinyIO_deploy increases slightly but remains lower than that of RoNIN ResNet, suggesting that the accuracy degradation due to reparameterization is within acceptable bounds.
+5(a), TinyIO effectively captures contextual motion information, yielding accurate and stable tracking.
+After reparameterization, the positioning error of TinyIO_deploy increases slightly but remains lower than that of R-ResNet, suggesting that the accuracy degradation due to reparameterization is within acceptable bounds.
 For complex trajectories (Fig.
-4(a)–(c)), TinyIO better models local fine-grained motion and exhibits substantially less drift than RoNIN ResNet when handling rotational or highly nonlinear motions; consequently, both TinyIO and TinyIO_deploy closely follow ground-truth trajectories, whereas RoNIN ResNet shows pronounced drift during turns and increasingly larger deviations as trajectory complexity grows.
-3.3 Ablation study
-We conduct ablation experiments to validate the reparameterization strategy and the contributions of individual modules (SGU and RepBlock).
-The middle section of Table 1 reports performance comparisons between TinyIO and TinyIO_deploy across the four datasets.
-After reparameterization, TinyIO_deploy (3.809M parameters) reduces the parameter count by approximately 14.52% relative to TinyIO (4.456M), and its model size is only 82.18% that of RoNIN ResNet (4.635M).
-Despite this compression, the increases in ATE and RTE for TinyIO_deploy are marginal, and TinyIO_deploy still outperforms RoNIN ResNet in localization accuracy.
+5(b), (c)), TinyIO better models local fine-grained motion and exhibits substantially less drift than R-ResNet when handling rotational or highly nonlinear motions; consequently, both TinyIO and TinyIO_deploy closely follow ground-truth trajectories, whereas R-ResNet shows pronounced drift during turns and increasingly larger deviations as trajectory complexity grows.
+IV-C Ablation study
+We conduct ablation experiments to validate the reparameterization strategy and the contributions of individual modules (DPAA and RepBlock).
+The Table I reports performance comparisons between TinyIO and TinyIO_deploy across the all datasets.
+After reparameterization, TinyIO_deploy (2.975M parameters) reduces the parameter count by approximately 16.67% relative to TinyIO (3.570M), and its model size is only 64.19% that of R-ResNet (4.635M).
+Despite this compression, the increases in ATE and RTE for TinyIO_deploy are marginal, and TinyIO_deploy still outperforms R-ResNet in localization accuracy.
 Visualization results in Fig.
@@
 These results indicate that the proposed reparameterization achieves a favorable trade-off between parameters and accuracy, making TinyIO particularly suitable for edge devices with limited storage.
-The right portion of Table 1 shows module ablation results.
-Removing SGU (TinyIO w/o SGU) leaves only the RepBlock: the model still captures local fine-grained motion and yields substantially lower localization error than RoNIN ResNet, indicating that RepBlock both reduces parameter complexity and contributes positively to accuracy.
-Integrating SGU into the full TinyIO architecture further improves accuracy, demonstrating that multi-scale feature fusion via SGU is critical for enhanced trajectory estimation.
+The Table II shows module ablation results.
+Removing DPAA (TinyIO (only RepBlock)) leaves only the RepBlock: the model still captures local fine-grained motion and yields substantially lower localization error than R-ResNet, indicating that RepBlock both reduces parameter complexity and contributes positively to accuracy.
+Integrating DPAA into the full TinyIO (RepBlock&DPAA) further improves accuracy, indicating that modeling contextual motion through DPAA is essential for enhancing trajectory estimation.
 These findings confirm the effectiveness and complementarity of the proposed components.
-4 Conclusion
+V Conclusion
 We present TinyIO, a lightweight IO framework that attains a favorable trade-off between localization accuracy and model parameters.
 In practical deployment, however, additional procedures such as model fine-tuning may be necessary to adapt to previously unseen carrier motion patterns.
-5 Acknowledgements
-This work is supported by Science and Technology Major Program of Fujian Province (No. 2022HZ026007) and partly supported by the Science and Technology Planning Project of Fujian province (2022I0001).